ydx+(1−3y)xy=3ydy
1
Expert's answer
2020-11-03T16:39:40-0500

"ydx+(1\u22123y)xy=3ydy"

A first order linear differential equation in the form of "y'(x)+p(x)y=q(x)"

"y+(1-3y)xy=3y\frac{dy}{dx}"

Substitute "\frac{dy}{dx}=y'"

"y+(1-3y)xy=3yy'"

Rewrite,

"y'+xy=\frac{1}{3}+\frac{x}{3}"

Integration Factor : "IF = e^{\frac{x^2}{2}}"

Differential Equation: "(IF\times y)'= IF\times q(x)"

"(e^{\frac{x^2}{2}} \times y)'= e^{\frac{x^2}{2}} \times \frac{1}{3} +e^{\frac{x^2}{2}} \times \frac{x}{3}"

"y = \frac{\sqrt 2 \sqrt \pi erfi \frac{x}{\sqrt 2}}{6e^{\frac{x^2}{2}}}+\frac{1}{3}+\frac {c_1}{e^{\frac{x^2}{2}}}"